# MiniBEE: A New Form Factor for Compact Bimanual Dexterity

### v2

@@ IV-B Experiments and Performance
We used the method above on a total of 3 tasks, each of them requiring bimanual dexterity for completion. All tasks are illustrated in Fig. 5.
@@ V Conclusion and Future Work
While MiniBEE exhibits a high level of bimanual dexterity even in a compact package, additional areas of improvement remain. In particular, in future work we aim to implement self-tracking of unimanual tasks executed relative to a fixed frame in the environment (possible in the case where the other gripper is kept fixed relative to the manipulation frame) in order to allow closed-loop policies for object acquisition from the environment. Other possible areas of improvement include better models of the MiniBEE position control system, to compensate for different gravity effects during data collection (grippers held by the operator) and policy execution (free-floating grippers). Nevertheless, we believe MiniBEE is a step towards highly dexterous bimanual systems easily deployed for a wide range of applications in human environments.

### v3

@@ IV-B Experiments and Performance
We used the method above on a total of 3 tasks, each of them requiring bimanual dexterity for completion. All tasks are illustrated in Fig. 5. For each task, there is a component where a single gripper must pick the object from a table in the environment. As stated, we are focusing only on demonstrating the intrinsic dexterity of our system, and therefore any pick up or placing of objects from the environment is a recorded trajectory. However, any manipulation between the grippers is done by a trained diffusion policy.
@@ V Conclusion and Future Work
While MiniBEE exhibits a high level of bimanual dexterity even in a compact package, additional areas of improvement remain. In particular, in future work we aim to implement self-tracking of uni-manual tasks executed relative to a fixed frame in the environment (possible in the case where the other gripper is kept fixed relative to the manipulation frame) in order to allow closed-loop policies for object acquisition from the environment. Other possible areas of improvement include better models of the MiniBEE position control system, to compensate for different gravity effects during data collection (grippers held by the operator) and policy execution (free-floating grippers). Finally, our KD metric used to assess different kinematics can be improved to include notions of torque, payload, and mass. Nevertheless, we believe MiniBEE is a novel approach to bimanual dexterity and a step towards highly dexterous bimanual systems that can easily deployed for a wide range of applications in human environments.